solve by elimination
x-3y=-16
-3x-3y=0

Found 2 solutions by checkley71, venugopalramana:
Answer by checkley71(8403)

X-3Y=-16 THEN -3Y=-16-X NOW SUBSTITUTINF -16-X FOR -3Y IN EQUATION #2 WE GET
-3X-16-X=0 OR -4X-16=0 OR -4X=16 OR X=16/-4 OR X=-4 THE -4-3Y=-16 OR -3Y=-12
OR Y=4

Answer by venugopalramana(3286)

solve by elimination
x-3y=-16............I
-3x-3y=0...........II
EQN.I-EQN.II
X+3X=16
4X=16
X=16/4=4
4-3Y=16
3Y=4-16=-12
Y=-4